# Next-Generation Database Interfaces: A Survey of LLM-based Text-to-SQL

### v3

@@ II-B2 Deep Learning-based Approaches
With the rise of deep neural networks, sequence-to-sequence models and encoder-decoder structures, such as LSTMs [79] and transformers [17], were adapted to generate SQL queries from natural language input [80, 19]. Typically, RYANSQL [19] introduced techniques like intermediate representations and sketch-based slot filling to handle complex questions and improve cross-domain generalization. Recently, researchers introduced graph neural networks (GNNs) for text-to-SQL tasks by leveraging schema dependency graphs to capture the relationships between database elements [18, 81].
@@ II-B3 PLM-based Implementation
Pre-trained language models (PLMs) have emerged as a powerful solution for text-to-SQL, leveraging the vast amounts of linguistic knowledge and semantic understanding captured during the pre-training process. The early adoption of PLMs in text-to-SQL primarily focused on fine-tuning off-the-shelf PLMs, such as BERT [24] and RoBERTa [82], on standard text-to-SQL datasets [13, 14]. These PLMs, pre-trained on large amounts of training corpus, captured rich semantic representations and language understanding capabilities. By fine-tuning them on text-to-SQL tasks, researchers aimed to leverage the semantic and linguistic understanding of PLMs to generate accurate SQL queries [80, 20, 83]. Another line of research focuses on incorporating schema information into PLMs to improve their understanding of database structures and enable them to generate more executable SQL queries. Schema-aware PLMs are designed to capture the relationships and constraints present in the database structure [21].
@@ II-B4 LLM-based Implementation
Large language models (LLMs), such as GPT series [84, 85, 86], have gained significant attention in recent years due to their capability to generate coherent and fluent text. Researchers have started exploring the potential of LLMs for text-to-SQL by leveraging their extensive knowledge reserves and superior generation capabilities [7, 9]. These approaches often involve prompt engineering to guide the proprietary LLMs in SQL generation [47] or fine-tuning the open-source LLMs on text-to-SQL datasets [9].
The integration of LLMs in text-to-SQL is still an emerging research area with significant potential for further exploration and improvement. Researchers are investigating approaches to better leverage the knowledge and reasoning capabilities of LLMs, incorporate domain-specific knowledge [33, 31], and develop more efficient fine-tuning strategies [10]. As the field continues to evolve, we anticipate the development of more advanced and superior LLM-based implementations that will elevate the performance and generalization capabilities of text-to-SQL to new heights.
@@ III Benchmarks & evalution
TABLE I: The statistics and analysis of well-known datasets of text-to-SQL ordered by release time. The original dataset indicates that the dataset is designed with a corresponding database, while post-annotated datasets involve annotating new components within existing datasets and databases rather than releasing a new database.
Original Dataset Release Time #Example #DB #Table/DB #Row/DB Characteristics
BIRD [33] May-2023 12,751 95 7.3 549K Cross-domain, Knowledge-augmented
KaggleDBQA [36] Jun-2021 272 8 2.3 280K Cross-domain
DuSQL [34] Nov-2020 23,797 200 4.1 - Cross-domain, Cross-lingual
SQUALL [44] Oct-2020 11,468 1,679 1 - Knowledge-augmented
CoSQL [35] Sep-2019 15,598 200 - - Cross-domain, Context-dependent
Spider [13] Sep-2018 10,181 200 5.1 2K Cross-domain
WikiSQL [14] Aug-2017 80,654 26,521 1 17 Cross-domain
Post-annotated Dataset Release Time Source Dataset Special Setting Characteristics
ADVETA [37] Dec-2022 Spider, etc. Adversarial table perturbation Robustness
Spider-SS&CG [38] May-2022 Spider Splitting example into sub-examples Context-dependent
Spider-DK [39] Sep-2021 Spider Adding domain knowledge Knowledge-augmented
Spider-SYN [40] Jun-2021 Spider Manual synonym replacement Robustness
Spider-Realistic [41] Oct-2020 Spider Removing column names in question Robustness
CSpider [42] Sep-2019 Spider Chinese version of Spider Cross-lingual
SParC [43] Jun-2019 Spider Annotate conversational contents Context-dependent
@@ III-A Datasets
As shown in Tab. I, we classify the datasets into ’Original Datasets’ and ’Post-annotated Datasets’ based on whether they were released with the original dataset and databases or created by adapting existing datasets and databases with special settings. For the original datasets, we provide a detailed analysis, including the number of examples, the number of databases, the number of tables per database, and the number of rows per database. For the post-annotated datasets, we identify their source dataset and describe the special setting applied to them. To illustrate the potential opportunities of each dataset, we annotate them based on their characteristics. The annotations are listed in the rightmost column of Tab. I, which we will discuss in detail below.
@@ III-A4 Robustness Dataset
Evaluating the accuracy of text-to-SQL systems with polluted or perturbed database contents (e.g., schema and tables) is crucial for assessing robustness. Spider-Realistic [41] removes explicit schema-related words from the NL questions, while Spider-SYN [40] replaces them with manually selected synonyms. ADVETA [37] introduces adversarial table perturbation (ATP), which perturbs tables by replacing original column names with misleading alternatives and inserting new columns with high semantic associations but low semantic equivalency. These perturbations lead to significant drops in accuracy, as a text-to-SQL system with low robustness may be misled by incorrect matches between tokens in NL questions and database entities.
@@ III-A5 Cross-lingual Dataset
SQL keywords, function names, table names, and column names are typically written in English, posing challenges for applications in other languages. CSpider [42] translates the Spider dataset into Chinese, identifying new challenges in word segmentation and cross-lingual matching between Chinese questions and English database contents. DuSQL [34] introduces a practical text-to-SQL dataset with Chinese questions and database contents provided in both English and Chinese.
@@ III-B1 Content Matching-based Metrics
SQL content matching metrics focus on comparing the predicted SQL query with the ground truth SQL query based on their structural and syntactic similarities.
@@ III-B2 Execution-based Metrics
  • Valid Efficiency Score (VES) [33] is defined to measure the efficiency of valid SQL queries. A valid SQL query is a predicted SQL query whose executed results exactly match the ground truth results. Specifically, VES evaluates both the efficiency and accuracy of predicted SQL queries. For a text dataset with $N$ examples, VES can be computed by:
@@
Most of the recent LLM-based text-to-SQL studies focus on these four datasets: Spider [13], Spider-Realistic [41], Spider-SYN [40], and BIRD [33]; and these three evaluation methods: EM, EX, and VES, we will primarily focus on them in the following analysis.
@@ IV Methods
The implementation of current LLM-based applications mostly relies on in-context learning (ICL) (prompt engineering) [87, 88, 89] and fine-tuning (FT) [90, 91] paradigms since the powerful proprietary models and well-architected open-source models are being released in large quantities [86, 45, 92, 93, 94, 95]. LLM-based text-to-SQL systems follow these paradigms for implementation. In this survey, we will discuss them accordingly.
@@ IV-A In-context Learning
Through extensive and widely recognized research, prompt engineering has been proven to play a decisive role in the performance of LLMs [96, 28], also impacting the SQL generation in different prompt styles [9, 46]. Necessarily, developing text-to-SQL methods in the in-context learning (ICL) paradigm is valuable for achieving promising improvement. The implementation of LLM-based text-to-SQL process to generate executable SQL query $Y$ can be formulated as:
$Y=f(\mathcal{Q},\mathcal{S},\mathcal{I}\mid\theta),$ (3)
@@
where $\mathcal{Q}$ represents the user question. $\mathcal{S}$ is the database schema/content, which can be decomposed as $\mathcal{S}=\langle\mathcal{C},\mathcal{T},\mathcal{K}\rangle$, where $\mathcal{C}=\{c_{1},c_{2},...\}$ and tables $\mathcal{T}=\{t_{1},t_{2},...\}$ represent the collection of various columns and tables, $\mathcal{K}$ is the potentially external knowledge (e.g. foreign key relationships [49], schema linking [30] and domain knowledge [33, 31]). $\mathcal{I}$ represents the instruction for the text-to-SQL task, which provides indicative guidance to trigger the LLMs for generating an accurate SQL query. $f(\cdot\mid\theta)$ is a LLM with parameter $\theta$. In the in-context learning (ICL) paradigm, we utilize an off-the-shelf text-to-SQL model (i.e., parameter $\theta$ of the model is frozen) for generating the predicted SQL query. Various well-designed methods within the ICL paradigm have been adopted for LLM-based text-to-SQL tasks. We group them into five categories $\mathbf{C}_{0:4}$, including $\mathbf{C}_{0}$-Trivial Prompt, $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement, the representative methods of each category are given in Tab. II.
$\mathbf{C}_{0}$-Trivial Prompt
@@
Trained through massive data, LLMs have a strong overall proficiency in different downstream tasks with zero-shot and few-shot prompting [97, 90, 98], which is widely recognized and used in real-world applications. In our survey, we categorized the above prompting approaches without the well-designed framework as trivial prompts (vanilla prompt engineering). As introduced above, Eq. 3 formulated the process of LLM-based text-to-SQL, which can also represent zero-shot prompting. The overall input $\mathcal{P}_{0}$ can be obtained by concatenating $\mathcal{I}$, $\mathcal{S}$ and $\mathcal{Q}$:
$\mathcal{P}_{0}=\mathcal{I}\oplus\mathcal{S}\oplus\mathcal{Q}.$ (4)
To regulate the prompting process, the OpenAI demonstration [^2] The prompt style that follows the official document from OpenAI platform: https://platform.openai.com/examples/default-sql-translate is set as the standard (trivial) prompt [30] for text-to-SQL.
TABLE II: Typical methods used for in-context learning (ICL) in LLM-based text-to-SQL. The full table of existing methods with categorization $\mathbf{C}_{1:4}$ and more details are listed in Tab. III.
@@
$\mathbf{C}_{0}$-Trivial Prompt Zero-shot [7] ChatGPT
Few-shot [9] ChatGPT
@@
TABLE III: Well-designed methods used in in-context learning (ICL) paradigm for LLM-based text-to-SQL ordered by release time. The methods are grouped in four categories based on their implementation perspective: $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, $\mathbf{C}_{4}$-Execution Refinement. The method in multiple categories will be introduced respectively. *There are multiple applied LLMs in the corresponding method; we present the selected LLM with representative performance. CoT method are reported in multiple venues: NeurIPS’23 [33], EMNLP’23 [51], VLDB’24 [9], arXiv’24 [32].
@@
SQL-PaLM [53] PaLM-2 [13] EX May-2023 arXiv’23
@@
FUSED [65] ChatGPT* [13] EX Feb-2024 arXiv’24
@@
Zero-shot: Many research works [7, 27, 46] utilize zero-shot prompting, studying mainly on the influence of the style of prompt construction and the zero-shot performance of various LLMs for text-to-SQL. As an empirical evaluation, [7] evaluates the baseline text-to-SQL capabilities of different early-developed LLMs [99, 85, 100] and the performance for different prompting styles. The results indicate that prompt design is critical for the performance, with error analysis, [7] propose more database content can harm the overall accuracy. Since ChatGPT emerged with impressive capabilities in conversational scenarios and code generation [101], [27] assesses its performance of text-to-SQL. With zero-shot settings, the results demonstrate that ChatGPT has a promising text-to-SQL performance compared to the state-of-the-art PLM-based systems. For fair comparability, [47] reveal effective prompt construction for LLM-based text-to-SQL; they study different styles of prompt construction and make conclusions of zero-shot prompt design based on the comparisons.
@@
Primary keys and foreign keys carry contiguous knowledge of different tables. [49] studies their impact by incorporating these keys into various prompt styles with different database content to analyze zero-shot prompting results. A benchmark evaluation [9] also studies the influence of foreign keys, with five different prompt representation styles, each style can be considered as the permutation and combination of the instruction, rule implication, and foreign key. Apart from the foreign key, this study also explores zero-shot prompting combined with the rule implication “no explanation” to collect concise outputs. Empowered by the annotated external knowledge of human experts, [33] follows the standard prompting and achieves improvement by incorporating the provided annotated oracle knowledge.
With the explosion of open-source LLMs, according to the results of similar evaluation, these models are also capable of zero-shot text-to-SQL task [45, 50, 46], especially code generation models [48, 46]. For zero-shot prompting optimization, [46] raises a challenge for designing an effective prompt template for LLMs; the former prompt construction lacks structure uniformity, which makes it hard to find out a concrete element within a prompt constructing template influences the performance of LLMs. They address this challenge by investigating a more unified series of prompt templates warping with different prefixes, infixes, and postfixes.
Few-shot: The technique of few-shot prompting is widely used in both practical applications and well-designed research, which has been proven efficient for eliciting better performance of LLMs [28, 102]. The overall input prompt of the few-shot approach LLM-based text-to-SQL can be formulated as an extension of Eq. 3:
$\mathcal{P}_{n}=\{\mathcal{F}_{1},\mathcal{F}_{2},\ldots,\mathcal{F}_{n}\}\oplus\mathcal{P}_{0},$ (5)
where the $\mathcal{P}_{n}$ represent the input prompt for $n$-shot learning, $n$ is the provided instances (examples) number; $\mathcal{F}$ denote the few-shot instance, which can be decomposed as $\mathcal{F}_{i}=(\mathcal{S}_{i},\mathcal{Q}_{i},Y_{i})$, $i$ is the serial number of instances. The study of few-shot prompting focuses on the number of representations and few-shot instance selection.
@@
As empirical studies, few-shot prompting for text-to-SQL are evaluated in multiple datasets with various LLMs [8, 32], exhibit a solid performance during the comparison with zero-shot prompting. [33] provides a 1-shot detailed example to trigger the text-to-SQL model for generating accurate SQL. [55] study the effect of the number of few-shot examples. [52] focus on the sampling strategies by studying the similarity and the diversity between different demonstrations, setting random sampling as the baseline, and evaluating different strategies and their combination for comparison. Furthermore, above the similarity-based selection, [9] evaluated masked question similarity selection and the upper limit of similarity approaches with various numbers of few-shot examples. A study of difficult-level samples selection [51] compared the performance of few-shot Codex [100], with random selection and difficulty-based selection for few-shot instances on difficulty categorized dataset [13, 41]. Three difficulty-based selection strategies are devised based on the number of selected samples at different difficulty levels. [49] utilize a hybrid strategy for selecting samples, which combines static examples and similarity-based dynamic examples for few-shot prompting. In their settings, they also evaluate the impact of different input schema styles and various static and dynamic exemplar numbers.

### v4

@@ II-B2 Deep Learning-based Approaches
With the rise of deep neural networks, sequence-to-sequence models and encoder-decoder structures, such as LSTMs [97] and transformers [15], were adapted to generate SQL queries from natural language input [98, 17]. Typically, RYANSQL [17] introduced techniques like intermediate representations and sketch-based slot filling to handle complex questions and improve cross-domain generalization. Recently, researchers introduced graph neural networks (GNNs) for text-to-SQL tasks by leveraging schema dependency graphs to capture the relationships between database elements [16, 99]. Despite their flexibility, these methods often generate incorrect or incomplete SQL syntax, such as missing clauses or improper nesting, due to limited structural awareness. They also struggle with rare or complex SQL operations, like nested subqueries and window functions, because such constructs are underrepresented in training data.
@@ II-B3 PLM-based Implementation
PLMs have emerged as a powerful solution for text-to-SQL, leveraging the vast amounts of linguistic knowledge and semantic understanding captured during the pre-training process. The early adoption of PLMs in text-to-SQL primarily focused on fine-tuning off-the-shelf PLMs, such as BERT [100] and RoBERTa [101], on standard text-to-SQL datasets [11, 12]. These PLMs, pre-trained on large amounts of training corpus, captured rich semantic representations and language understanding capabilities. By fine-tuning them on text-to-SQL tasks, researchers aimed to leverage the semantic and linguistic understanding of PLMs to generate accurate SQL queries [98, 18, 102]. Another line of research focuses on incorporating schema information into PLMs to improve their understanding of database structures and enable them to generate more executable SQL queries. Schema-aware PLMs are designed to capture the relationships and constraints present in the database structure [19]. While PLMs generate more accurate SQL than deep learning-based methods, they still face difficulties with complex operations like outer joins or aggregations, often producing syntactically flawed queries. Besides that, cross-domain performance drops significantly when schema structures or vocabulary differ from training data, necessitating resource-intensive fine-tuning.
@@ II-B4 LLM-based Implementation
Large language models (LLMs), such as the GPT series [103, 104, 105], have gained significant attention in recent years due to their ability to generate coherent and fluent text. Researchers have begun exploring the potential of LLMs for text-to-SQL by leveraging their extensive knowledge reserves and superior generation capabilities [6, 8]. These approaches often involve prompt engineering to guide proprietary LLMs in SQL generation [106] or fine-tuning open-source LLMs on text-to-SQL datasets [8].
The integration of LLMs in text-to-SQL is still an emerging research area with significant potential for further exploration and improvement. Researchers are investigating ways to better leverage the knowledge and reasoning capabilities of LLMs, incorporate domain-specific knowledge [1, 69], and develop more efficient fine-tuning strategies [24]. As the field continues to evolve, we anticipate the development of more advanced and superior LLM-based implementations that will elevate the performance and generalization of text-to-SQL to new heights.
@@ III Benchmarks & evalution
TABLE I: The statistics and analysis of well-known datasets of text-to-SQL ordered by release time. The original dataset indicates that the dataset is designed with a corresponding database, while post-annotated datasets involve annotating new components within existing datasets and databases rather than releasing a new database. For each dataset, the languages utilized are denoted by the following abbreviations: EN for English, ZH for Chinese, and VI for Vietnamese.
Original Dataset Release Time #Example #DB #Table/DB #Row/DB Language Characteristics
BIRD-CRITIC [1] Feb-2025 600 - - - EN Knowledge-augmented, Long-context
Spider 2.0 [42] Nov-2024 632 213 - - EN Knowledge-augmented, Long-context
BULL [45] Jan-2024 4,966 3 26 - EN Specialized-domain
BIRD [1] May-2023 12,751 95 7.3 549K EN Cross-domain, Knowledge-augmented
KaggleDBQA [33] Jun-2021 272 8 2.3 280K EN Cross-domain
DuSQL [31] Nov-2020 23,797 200 4.1 - EN Cross-domain, Cross-lingual
SQUALL [41] Oct-2020 11,468 1,679 1 - EN Knowledge-augmented
CoSQL [32] Sep-2019 15,598 200 - - EN Cross-domain, Context-dependent
Spider [11] Sep-2018 10,181 200 5.1 2K EN Cross-domain
WikiSQL [12] Aug-2017 80,654 26,521 1 17 EN Cross-domain
Post-annotated Dataset Release Time Source Dataset Special Setting Language Characteristics
Dr. Spider [44] Jan-2023 Spider Perturbations in DB, query and SQL EN Robustness
ADVETA [34] Dec-2022 Spider, etc. Adversarial table perturbation EN Robustness
Spider-SS&CG [35] May-2022 Spider Splitting example into sub-examples EN Context-dependent
Spider-DK [36] Sep-2021 Spider Adding domain knowledge EN Knowledge-augmented
Spider-SYN [37] Jun-2021 Spider Manual synonym replacement EN Robustness
Spider-Vietnamese [43] Nov-2020 Spider Vietnamese version of Spider VI Cross-lingual
Spider-Realistic [38] Oct-2020 Spider Removing column names in question EN Robustness
CSpider [39] Sep-2019 Spider Chinese version of Spider ZH Cross-lingual
SParC [40] Jun-2019 Spider Annotate conversational contents EN Context-dependent
@@ III-A Datasets
As shown in Tab. I, we categorize the datasets into “Original Datasets” and “Post-annotated Datasets” based on whether they were released with the original dataset and databases or developed by adapting existing datasets and databases with special settings. For the original datasets, we offer a detailed analysis, including the number of examples, databases, tables per database, and rows per database. For post-annotated datasets, we identify their source dataset and describe the special settings applied to them. To highlight the potential opportunities of each dataset, we annotate them according to their characteristics. The annotations are listed in the rightmost column of Tab. I, which we will discuss in detail below.
@@ III-A4 Robustness Dataset
Evaluating the accuracy of text-to-SQL systems with polluted or perturbed database contents (e.g., schema and tables) is crucial for assessing robustness. Spider-Realistic [38] removes explicit schema-related words from the NL questions, while Spider-SYN [37] replaces them with manually selected synonyms. ADVETA [34] introduces adversarial table perturbation (ATP), which perturbs tables by replacing original column names with misleading alternatives and inserting new columns with high semantic associations but low semantic equivalency. Based on the generation ability of large-scale pre-trained language models, Dr. Spider [44] provided 17 different perturbation strategies on databases, natural language queries and SQL statements to measure the robustness of text-to-SQL system. These perturbations lead to significant drops in accuracy, as a text-to-SQL system with low robustness may be misled by incorrect matches between tokens in NL questions and database entities.
@@ III-A5 Cross-lingual Dataset
The SQL keywords, function names, table and column names are typically written in English, posing challenges for applications in other languages. CSpider [39] addresses this by translating the Spider dataset into Chinese, identifying new challenges in word segmentation and cross-lingual matching between Chinese questions and English database contents. DuSQL [31] introduces a practical text-to-SQL dataset with Chinese questions and database contents provided in both English and Chinese. Spider-Vietnamese [43] translates the Spider dataset into Vietnamese, empirically revealing several strategies to enhance text-to-SQL methods within the Vietnamese context.
III-A6 Long-context dataset
In real-world database applications, numerous challenging scenarios demand complex SQL queries with significant token lengths, often exceeding 100 tokens, involving multiple intricate operations. Spider 2.0 [42] exemplifies this complexity by presenting enterprise-level text-to-SQL problems that require advanced reasoning across various SQL queries and dialects. This benchmark reveals the limitations of current models, as demonstrated by lower success rates—with advanced LLMs solving only 17.1% of tasks—underscoring its unique challenges. In parallel, BIRD-CRITIC [^2] https://bird-critic.github.io/ [1] introduces a diagnostic dimension, comprising 600 tasks designed to assess the efficacy of LLMs in resolving real-world user issues across diverse SQL dialects. Together, these datasets push the boundaries of text-to-SQL research, emphasizing the challenges of complex, long-context SQL environments.
III-A7 Specialized-domain dataset
Despite advancements in text-to-SQL research, there is still a shortage of specialized datasets tailored for domain-specific analysis. To address this gap, BULL [45] introduces a dedicated text-to-SQL benchmark for the financial sector. The dataset encompasses databases related to funds, stocks, and macroeconomic data, offering a comprehensive platform to evaluate and enhance text-to-SQL methods in financial applications, effectively tackling the financial domain unique text-to-SQL challenges.
@@ III-B1 Content Matching-based Metrics
SQL content matching metrics focus on comparing the predicted SQL with the ground truth based on their structural and syntactic similarities.
@@ III-B2 Execution-based Metrics
  • Valid Efficiency Score (VES) [1] is defined to measure the efficiency of valid SQL queries. A valid SQL query is a predicted SQL whose executed results exactly match the ground truth results. Specifically, VES evaluates both the efficiency and accuracy of predicted SQL queries. For a text dataset with $N$ examples, VES is computed by:
@@
Most of the recent LLM-based text-to-SQL studies focus on BIRD and Spider-related datasets and execution-based metrics, the corresponding details can be found at Tab. III.
@@ IV Methods
The implementation of the most current LLM-based text-to-SQL methods relies on in-context learning (ICL) [107, 108, 109] and fine-tuning (FT) [110, 111] paradigms due to the release of powerful proprietary and well-architected open-source LLMs in large quantities [105, 112, 113, 114, 115, 116]. In this section, we will discuss these paradigms with detailed categorization accordingly.
@@ IV-A In-context Learning
Through extensive and widely recognized research, ICL (also called prompt engineering) has been proven to play a decisive role in the performance of LLMs across various tasks [117, 26, 118]. Consequently, it also impacts SQL generation across different prompt styles [8, 119]. The ICL implementation of the LLM-based text-to-SQL process to generate a predicted SQL query $\hat{Y}$, can be formulated as:
$\hat{Y}=\pi\left(I,Q,\mathcal{S}\mid\theta\right),$ (3)
@@
where $I$ represents the instruction for the text-to-SQL task, $Q$ represents the user question. $\mathcal{S}$ is the database schema/content, which can be decomposed as $\mathcal{S}=\langle\mathcal{C},\mathcal{T},\mathcal{K}\rangle$. Here, $\mathcal{C}=\{c_{1},c_{2},...\}$ and $\mathcal{T}=\{t_{1},t_{2},...\}$ represent the collection of various columns $c$ and tables $t$, respectively. $\mathcal{K}$ is the potentially supplementary knowledge, such as foreign key relationships [67], schema linking [28, 29], and external knowledge [1, 69]. $\pi(\cdot\mid\theta)$ is a LLM $\pi$ with parameter $\theta$. In the ICL paradigm, LLM-based methods primarily utilize an off-the-shelf text-to-SQL model, meaning the parameter $\theta$ of the model is frozen, for generating the predicted SQL query.
@@
We further specify the prompt engineering strategies in the ICL paradigm as zero-shot and few-shot prompting. Trained through massive data, LLMs have a strong overall proficiency in different downstream tasks with zero-shot and few-shot prompting [120, 110, 121], which is widely recognized and utilized in real-world applications. Formally, the overall input $P_{0}$ of zero-shot prompting for text-to-SQL can be obtained by concatenating $I$, $\mathcal{S}$, and $Q$ as introduced above:
$P_{0}=I\oplus\mathcal{S}\oplus Q.$ (4)
Subsequently, the overall input $P_{k}$ for few-shot prompting (specifically, $k$-shot) can be formulated as an extension of $P_{0}$:
$P_{k}=\{\mathcal{E}_{1},\mathcal{E}_{2},\ldots,\mathcal{E}_{k}\}\oplus P_{0},$ (5)
where $k$ is the number of provided instances (examples) in few-shot prompting. $\mathcal{E}_{i}$ represents the $i$-th few-shot instance, each of which can be decomposed as $\mathcal{E}_{i}=\left(\mathcal{S}_{i},Q_{i},Y_{i}\right)$, where $Y_{i}$ is the gold SQL query of $Q_{i}$. The index $i\in\{1,\ldots,k\}$ also serves as the corresponding serial number for $\mathcal{S}_{i}$, $Q_{i}$, and $Y_{i}$. The OpenAI demonstration [^3] The prompt format follows the official document provided by the OpenAI platform: https://platform.openai.com/examples/default-sql-translate provides an indicative example and regulation for the prompting process in LLM-based text-to-SQL [28]. The prompt designs and formats utilized by a substantial portion of the ICL methods discussed in this survey are developed based on the OpenAI regulation. Furthermore, most of these ICL methods incorporate zero-shot and few-shot prompting into their text-to-SQL frameworks.
As the mainstream research approach in LLM-based text-to-SQL, various well-designed methods within the ICL paradigm have been adopted to enhance LLM’s SQL generation. In this section, we group these existing methods into five categories, $\mathbf{C}_{0:4}$, including $\mathbf{C}_{0}$-Vanilla Prompting, $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement. The representative methods of each category are provided in Tab. II. We will introduce each category accordingly and summarize their characteristics at the beginning of each corresponding paragraph.
IV-A1 $\mathbf{C}_{0}$-Vanilla Prompting
In our survey, we identified a category of prompting strategies that do not propose a well-designed ICL framework. These strategies are based solely on zero-shot and few-shot prompting. Despite their simplicity, they offer valuable insights and findings on the ICL paradigm in LLM-based text-to-SQL. We refer to these strategies as vanilla prompting strategies and further divide their studies into zero-shot and few-shot studies.
TABLE II: A summary of representative methods used in the ICL paradigm for LLM-based text-to-SQL is provided. The complete table, along with a detailed categorization of existing well-designed methods denoted as $\mathbf{C}_{1:4}$, is available in Tab. III. *Note that SimpleDDL is a prompting style that provides only table and column names [119], rather than a method.
@@ IV-A1 $\mathbf{C}_{0}$-Vanilla Prompting
$\mathbf{C}_{0}$-Vanilla Prompting SimpleDDL [119]* SQLCoder
@@
TABLE III: Well-designed methods in the ICL paradigm for LLM-based text-to-SQL are ordered by release time. The methods are grouped into four categories based on their implementation perspective: $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement. Methods that fall into multiple categories will be introduced accordingly. The dataset column consists of six text-to-SQL datasets, which can be clarified by names: Spider [11], BIRD [1], Spider-DK [36] Spider-SYN [37], Spider-Realistic [38], and Spider 2.0 [42]. We use the reference ID in the table due to column space limitations. *There are multiple applied LLMs in the corresponding method; the model with representative performance is presented. CoT method is introduced in multiple venues: NeurIPS’23 [1], EMNLP’23 [47], VLDB’24 [8], arXiv’24 [30].
@@
SQL-PaLM [73] PaLM-2 [11] EX May-2023 TMLR’24
@@
FUSED [68] ChatGPT* [11] EX Feb-2024 EMNLP’24 Findings
@@
Dubo-SQL [70] ChatGPT* [1] EX Apr-2024 arXiv’24
TA-SQL [70] GPT-4* [11, 1, 36, 38] EX Apr-2024 ACL’24 Findings
MCS-SQL [55] GPT-4 [11, 1] EX, VES May-2024 COLING’25
CHESS [56] GPT-4o* [11, 1, 42] EX, VES May-2024 arXiv’24
SuperSQL [57] GPT-4* [11, 1] EX, EM, VES Jun-2024 VLDB’24
Distillery [58] GPT-4o-Mini* [1] EX Aug-2024 NeurIP’24 Workshop
E-SQL [80] GPT-4o* [11, 1] EX, VES Sep-2024 arXiv’24
CHASE-SQL [76] Gemini-1.5* [11, 1] EX, VES Oct-2024 ICLR’25
RSL-SQL [59] DeepSeek-Chat* [11, 1] EX, VES Nov-2024 arXiv’24
ICRL [71] ChatGPT [11] EX Jan-2025 COLING’25
Gen-SQL [60] LLaMA-3* [11, 1] EX, EM, VES Jan-2025 COLING’25
SAFE-SQL [72] GPT-4o-Mini* [11] EX, EM Feb-2025 arXiv’25
ReFoRCE [61] o1-preview* [42] EX Feb-2025 arXiv’25
@@
Zero-shot studies: The zero-shot-based ICL strategies primarily focus on studying the influence of prompt representation styles and the baseline performance of various LLMs for text-to-SQL [6, 25, 119]. As an empirical evaluation, [6] evaluates the baseline text-to-SQL capabilities of various early-developed LLMs [122, 104, 123] and their performance across different prompting styles. The experimental results indicate that prompt design is critical for performance. Through error analysis, this study [6] further suggests that including more database content can harm the overall accuracy of text-to-SQL. Since the emergence of ChatGPT [104] with its impressive capabilities in conversational scenarios and code generation [124], [25] evaluates its performance on Spider [11] and its variant [37, 36, 38] datasets. Under zero-shot settings, the results demonstrate that ChatGPT has promising text-to-SQL performance compared to state-of-the-art PLM-based systems. Aside from the proprietary models discussed above, similar evaluations indicate that open-source LLMs can also perform zero-shot text-to-SQL tasks [112, 82, 119], particularly code-specific models [79, 119].
@@
Primary keys and foreign keys convey continuous knowledge across different tables, making them an effective prompting style in zero-shot scenarios. [67] studied their impact by incorporating these keys into various prompt representation styles with different database content to analyze performance outcomes. A benchmark evaluation [8] also assessed the influence of foreign keys using five distinct prompt representation styles. Each style can be viewed as a permutation and combination of instruction, rule implication, and foreign key elements. In addition to using foreign keys, this study explores zero-shot prompting combined with the rule implication “no explanation” to produce concise outputs. With the help of annotated external knowledge from human experts, the evaluation in the BIRD benchmark [1] significantly improves by incorporating the provided annotated oracle knowledge.
In terms of zero-shot prompting optimization, [119] highlights the challenge of designing effective prompt templates for LLMs. Previous simple prompt constructions often lack structural uniformity, complicating the identification of specific elements in a prompt template that influence LLM performance. This challenge is addressed by investigating a more unified set of prompt templates that include various prefixes, infixes, and postfixes. To ensure fair comparability, [106] explores effective prompt construction for LLMs by studying different styles of prompt representation and highlights their findings on the crucial role of table structure, content, prompt length, and optimal representations in effective prompting.
@@
Few-shot studies: The study of few-shot prompting strategies focuses on the number and selection of few-shot instances. Empirical studies have evaluated few-shot prompting for text-to-SQL across multiple datasets using various LLMs [7, 30], demonstrating solid improvements compared to zero-shot prompting. [1] provides a detailed 1-shot example to trigger the text-to-SQL model for generating accurate SQL. [74] explores the effect of the number of few-shot examples. [64] focuses on sampling strategies by examining the similarity and diversity between different demonstrations, setting random sampling as the baseline, and evaluating various strategies and their combinations for comparison. Furthermore, beyond similarity-based selection, [8] evaluates masked question similarity selection and the upper limit of similarity approaches with various numbers of few-shot examples.
A study on difficulty-level sample selection [47] compared the performance of few-shot Codex [123], using random selection and difficulty-based selection for few-shot instances on difficulty-categorized datasets [11, 38]. Three difficulty-based selection strategies are devised based on samples selected by difficulty level. [67] utilizes a hybrid strategy for selecting samples, combining static examples and similarity-based dynamic examples for few-shot prompting. In their settings, they also evaluate the impact of different input schema styles and various numbers of static and dynamic exemplars.